OpenStudy (anonymous):

The equation - 5x² + 72x + 378 models the number of students enrolled in a school where x is the number of years since the school first opened in 1990. In which year were 598 students enrolled?

OpenStudy (anonymous):

-5x^2+72x+378=598

OpenStudy (nurali):

5x² + 72x + 378 = 598 -5x² + 72x = 220 5x² - 72x + 220 = 0 Solving by quadratic formula, x = 4.4 years or 10 years

OpenStudy (anonymous):

The available answers are 1992, 1998, 2000, 2010

OpenStudy (nurali):

1990+10=2000

OpenStudy (anonymous):

Thanks!!

OpenStudy (nurali):

My Pleasure.